A uniform board that is 1.50m long and weighs 3.50x10^2 N is suspended horizontally by two vertical cables at each end. Cable A can support a maximum tension of 5.00x10^2 N without breaking, and cable B can support up to 4.00x10^2 N. You want to place a small weight on this bar. Find the heaviest weight you can put on without breaking either cable and its location.


1
Expert's answer
2021-07-09T08:36:00-0400

One of the condition of equilibrium is that sum of upward force must be equal to sum of downward force

"T_A+T_B=W+350\\
T_A=500N,T_B=400N\\
500+400=W+350\\
W=900-350=550N\\"

Taking moment about the end cable A is attached

"W\times{x}+350\times{0.75}=T_B\times{1.5}\\
550\times{x}+262.5=600\\
550\times{x}=600-262.5\\
 550\times{x}=337.5\\
x=\dfrac{337.5}{550}\\
x=0.614m"

The heaviest weight is 550N placed at 0.614m from cable A